A compound has an empirical formula of C7H12O and a formula mass of 224 g/mol. What is the compound’s molecular formula?


1
Expert's answer
2021-08-11T05:22:10-0400

Empirical formula = "C_7H_{12}O"

Mass of the empirical formula compound = "((12\u00d77)+(1\u00d712)+(16))="

"84+12+16 = 112"g/mol


(Empirical formula)n = Molecular formula

"\therefore" n × Empirical mass = Molecular (Formula) mass

n × 112 = 224

n = 2


Molecular formula = "(C_7H_{12}O)_n"

"= (C_7H_{12}O)_2 = C_{14}H_{24}O_2"